Oct 31, 2019 Calorimeter

From GlueXWiki
Jump to: navigation, search

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book

Goals for Calorimetry Group

  1. Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
  2. Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at east 5 MeV.
  3. Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.

Action Items

Short term items:

  1. Summarize spring 2019 gain calibrations before uploading to ccdb (Karthik)
  2. Take a new noise measurements of the BCAL FADC baselines (Mark)
  3. Complete Calorimeter work packages

Long term items:

  1. From Mike's work: -cos(phi) type dependence in the data and a phi dependence at the downstream end of the bcal. See Sep 21, 2017 Calorimeter Meeting.
  2. FCAL: It has been noticed that the pi0 mass shifts when the LED pulser runs at high rate. This is still not understood. See Log Entry 3595928 and references. Several runs have been taken in raw mode. These runs exhibit the shift in pi0 mass with LED running at 1 kHz. There is a need to analyze these to try to determine the cause:
    1. 50634 - FCAL Green LED: 1kHz, 29 V, Livetime is 9%, 1.7 M events
    2. 50635 - FCAL Violet LED: 1kHz, 22 V, Livetime is 9%, 1.4 M events
    3. 50636 - FCAL Blue LED: 1kHz, 15 V, Livetime is 9%, 1.5 M events

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Run Updates
    1. FCAL
    2. BCAL
  4. Calibrations
    1. Nonlinear corrections File:Spring 2019 nonlinearity.pdf
  5. Monitoring
  6. Efficiencies
  7. Simulations
    1. Tolga: Not simulation actually, but neutral ID project related. Data first look at pi+pi-pi0 for photon timing dists in BCAL. [2]
  8. NIM articles
  9. Any other business


Attending: Elton, Mark, Colin, Jon (JLab); Will (CMU); Churamani, Tolga (FIU); Christina, George (Athens); Matt (IU); Sean (FSU); Zisis, Ahmed, Varun, Karthik (UofR).

  1. Announcements
  2. Action Items
    • See reports
  3. Run Updates
    1. FCAL (Colin)
      1. About 30 bases need replacing. Nick is working on spares
      2. Sasha installed the new firmware on half of the FCAL. No issues found yet, but comprehensive testing still needed.
      3. Four FADC boards are not showing hits. (Appears unrelated to new firmware). Sasha will check and replace if needed. Elton: Could check pedestals.
      4. (Elton) There are tests planned using the GEM (10x10 cm2) located immediately upstream of DIRC for test runs.
        • Mark had suggested that these data (with straight tracks) might be useful for calibrating the position resolution of the FCAL.
        • This is a small area, but might provide improved position resolution at the FCAL. The effects of multiple scattering have not yet been evaluated.
    2. BCAL (Jon)
      1. Jon is learning how to check the BCAL
        • Turned everything on.
        • Chillers are now operational, but there is a read back problem with the chiller temperature for US chiller.
        • Jon took LED data according to indications from Ahmed. At his request, Jon will move files to work disk so that Ahmed can have easy access.
        • New voltage file has been created. Jon still needs to compare to last year to check there are minor changes.
  4. Calibrations
    1. Nonlinear corrections File:Spring 2019 nonlinearity.pdf (Karthik)
      • Karthik presented the analysis of the non-linearity corrections for the 2019 data.
      • Attempts were made to extrapolate to higher energy using the single-shower energies using asymmetric decays.
      • Elton: Proposal to obtain the non-linearity corrections for the 2019 data by using the same procedures that were used or 2018.
      • New procedures to extrapolate to higher energies should be presented in a little more detail next time to inform the decision on whether to change procedures in the future.
  5. Monitoring
  6. Efficiencies
  7. Simulations
    1. Tolga: Not simulation actually, but neutral ID project related. Data first look at pi+pi-pi0 for photon timing dists in BCAL. [3]
      • Uses photons from data in the reaction gp -> pi+pi-pi0p, pi0->gg, to study the timing of the photons in the BCAL. He finds some offsets in both the velocity (beta) and timing relative to the RF (~0.2 ns).
      • Mark: Does not expect these offsets in the data. Check the monitoring_hists to see if it is also seen there.
  8. NIM articles
    1. Matt will take a look at the existing draft and give feedback
    2. Mark has studied the width of etas in gp -> pi+pi-eta, eta->gg to provide a performance plot for BCAL and FCAL in the NIM paper. Finds reasonable agreement for BCAL, but width of FCAL is 20-30% lower in MC than in data. We believe that mcsmear needs to be tuned to reproduced the widths observed in data. It appears that this discrepancy has existed for some time.
    3. The comparison plots can be found in links to yesterday's Production/Analysis meeting
  9. Any other business